OpenStudy (dj3strella):

Help me please 8d^3/6c divided by c^3d^4/24

OpenStudy (dj3strella):

c^2d^7/6c 16c^2/d 64c^2/d 32/c^4d

OpenStudy (jdoe0001):

\(\bf \cfrac{ \frac{8d^3}{6c} }{ \frac{c^3d^4}{24} }\qquad recall\implies {\color{brown}{ \cfrac{\quad \frac{a}{b}\quad }{\frac{c}{d}}\implies \cfrac{a}{b}\cdot \cfrac{d}{c}}}\qquad thus \\ \quad \\ \cfrac{ \frac{8d^3}{6c} }{ \frac{c^3d^4}{24} }\implies \cfrac{8d^3}{6c}\cdot \cfrac{24}{c^3d^4}\)

OpenStudy (dj3strella):

so then i multiply

OpenStudy (jdoe0001):

yeap\(\bf \cfrac{8d^3}{6c}\cdot \cfrac{24}{c^3d^4}\implies \cfrac{8\cancel{ d^3 }\cdot \cancel{ 24 }}{\cancel{ 6c }\cdot c^3\cancel{ d^4 }}\implies ?\)

OpenStudy (jdoe0001):

hmmm actually \(\bf \cfrac{8d^3}{6c}\cdot \cfrac{24}{c^3d^4}\implies \cfrac{8\cancel{ d^3 }\cdot \cancel{ 24 }}{\cancel{ 6 }c\cdot c^3\cancel{ d^4 }}\implies ?\)

OpenStudy (dj3strella):

8/c^4

OpenStudy (jdoe0001):

\(\bf \cfrac{8d^3}{6c}\cdot \cfrac{24}{c^3d^4}\implies \cfrac{8d^3}{6c}\cdot \cfrac{6\cdot 4}{c^3d^3\cdot d}\implies \cfrac{8\cancel{ d^3 }\cdot \cancel{ 6}\cdot 4 }{\cancel{ 6 }c\cdot c^3\cancel{ d^3 }\cdot d}\implies ?\)

OpenStudy (dj3strella):

so its 32/c^4d

OpenStudy (jdoe0001):

yeap

OpenStudy (dj3strella):

thanks can you help me with something else

OpenStudy (jdoe0001):

sure just post anew, so if I dunno someone else may and we can also revvise each other

OpenStudy (dj3strella):

ok